# A Survey on 3D Gaussian Splatting

### v2

@@ 3.2.1 Parameter Optimization
$\bullet$ Loss Function. Once the synthesis of the image is completed, the difference between the rendered image and ground truth can be measured. All learnable parameters are optimized by stochastic gradient descent using the L1 and D-SSIM loss functions:
@@
where $\lambda$ is a weighting factor. The loss function of 3D GS is slightly different from that of NeRFs. NeRFs typically calculate losses at the pixel level rather than the image level because of the costly ray-marching.
@@ 3.2.2 Density Control
$\bullet$ Initialization. 3D GS starts with the initial set of sparse points from SfM or random initialization. Afterwards, point densification and pruning are adopted to control the density of 3D Gaussians.
$\bullet$ Point Densification. In the point densification phase, 3D GS adaptively increases the density of Gaussians to better capture the details of a scene. This process particularly focuses on areas with missing geometric features or regions where Gaussians are too spread out. The densification procedure will be performed at regular intervals (i.e., after a certain number of training iterations), focusing on those Gaussians with large view-space positional gradients (i.e., above a specific threshold). It involves either cloning small Gaussians in under-reconstructed areas or splitting large Gaussians in over-reconstructed regions. For cloning, a copy of the Gaussian is created and moved towards the positional gradient. For splitting, a large Gaussian is replaced with two smaller ones, reducing their scale by a specific factor. This step seeks an optimal distribution and representation of Gaussians in the 3D space, enhancing the overall quality of the reconstruction.
$\bullet$ Point Pruning. The point pruning stage involves the removal of superfluous or less impactful Gaussians, which can be viewed as a regularization process to some extent. This step is executed by eliminating Gaussians that are virtually transparent (with $\alpha$ below a specified threshold) and those that are excessively large in either world-space or view-space. In addition, to prevent unjustified increases in Gaussian density near input cameras, the alpha value of the Gaussians is set close to zero after a certain number of iterations. This allows for a controlled increase in the density of necessary Gaussians while enabling the culling of redundant ones. The process not only helps in conserving computational resources but also ensures that the Gaussians in the model remain precise and effective for the representation of the scene.
@@
The last months have witnessed significant advancements in the realm of 3D GS. It brings us great pleasure to witness the fruition of the research directions we noted in our previous editions. In the subsequent sections, we seek to elaborate on select extended versions of 3D GS. These are: i) Data-efficient 3D GS [52, 53, 54, 55, 56, 57, 58, 59, 60, 61] (Sec. 4.1), ii) Memory-efficient 3D GS [62, 63, 64, 65, 66, 67, 68] (Sec. 4.2), iii) Photorealistic 3D GS [69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85] (Sec. 4.3), iv) Improved Optimization Algorithms [86, 87, 82, 88, 89, 90] (Sec. 4.4), v) 3D Gaussian with More Properties [91, 92, 93, 94, 95, 96, 97] (Sec. 4.5), and vi) 3D GS with Structured Information [98, 99, 100] (Sec. 4.6).
@@ 4.1 Data-efficient 3D GS
A notable issue of 3D GS is the emergence of artifacts in areas with insufficient observational data. This challenge is a prevalent limitation in radiance field rendering, where sparse data often leads to inaccuracies in reconstruction. From a practical perspective, reconstructing scenes from limited viewpoints is of significant interest, particularly for their potential to enhance realism and user experience with minimal input.
Two primary strategies are used for data-efficient 3D GS. i) The first introduces additional constraints such as depth information to enhance the detail and global consistency. For example, DNGaussian [56] introduced a depth-regularized approach to address the challenge of geometry degradation in sparse input views through hard and soft depth regularization and global-local depth normalization. It rectified the geometry using depth information extracted from pre-trained monocular depth estimators and optimized the positioning of Gaussian primitives without changing their shape, ensuring a balance between detailed color appearance and coherent scene geometry. ii) The second generates 3D Gaussians that can be used for rendering directly without optimization, by using deep neural networks. This paradigm typically requires multiple views for training but can reconstruct the 3D scene with only one input image. For instance, PixelSplat [54] proposed to sample Gaussians from dense probability distributions. It incorporated a multi-view epipolar transformer and a reparameterization trick to avoid local minima and maintain gradient flow. Splatter Image [55] applied GS in a monocular setting through a learning-based approach, utilizing a 2D image-to-image network that maps an input image to a 3D Gaussian per pixel. Gamba [61] introduced an end-to-end, Mamba based sequential network [101] to generate 3D Gaussians with linear scalability. Note that this paradigm is mainly focused on the reconstruction of objects, and its generalizability leaves a lot of room for improvement.
@@ 4.2 Memory-efficient 3D GS
While 3D GS demonstrates remarkable capabilities, its scalability poses significant challenges, particularly when juxtaposed with NeRF-based methods. The latter benefits from the simplicity of storing merely the parameters of a learned MLP. This scalability issue becomes increasingly acute in the context of large-scale scene management, where the computational and memory demands escalate substantially. Consequently, there is an urgent need to optimize memory utilization during both the training phase and the storage of the model.
There are two main directions to reduce memory usage. i) The first involves reducing the number of 3D Gaussians, i.e., pruning insignificant 3D Gaussians. For example, LightGaussian [62] pruned insignificant Gaussians based on their contribution to scene reconstruction. Lee et al. [64] introduced a new volume-based masking strategy that efficiently reduces the quantity of Gaussians required for scene representation without compromising performance. ii) The second category focuses on compressing the memory usage of 3D Gaussian properties. For instance, LightGaussian [62] introduced vector quantization [102] for attribute quantization to achieve lower bitwidth representations. Lee et al. [64] employed a grid-based neural field for a more compact representation of view-dependent colors and utilizes a codebook approach for the concise encoding of Gaussian geometries. Niedermayr et al. [67] compressed color and Gaussian parameters into compact codebooks, using sensitivity measures for effective quantization and fine-tuning, coupled with entropy encoding to utilize spatial coherence.
@@ 4.3 Photorealistic 3D GS
The current rendering pipeline of 3D GS (Sec. 3.1) is straightforward and involves several drawbacks. For instance, the simple visibility algorithm may lead to a drastic switch in the depth/blending order of Gaussians. The realness of rendered images (e.g., aliasing, reflections, and artifacts) can be further optimized.
Here we list several key points for enhancing realness. i) Varying Resolutions. Due to the discrete sampling paradigm (viewing each pixel as a single point instead of an area), 3D GS is susceptible to aliasing when dealing with varying resolutions, which leads to blurring or jagged edges. Yan et al. [71] argued that this is mainly because traditional rendering methods cannot effectively manage the disparity between the sampling frequency of pixels and the high-frequency details of the scene, leading to visual artifacts and performance issues. Hence, they introduced multi-scale 3D GS, where the scene is represented using Gaussians of varying sizes. Analytic-Splatting [83] adopted an analytic approximation of the Gaussian integral within the pixel area, leveraging a conditioned logistic function for the cumulative distribution function to better capture the intensity response of pixels. It aimed to make the rendering process sensitive to changes in pixel footprints at different resolutions, thereby enhancing anti-aliasing capabilities and preserving high-quality details. ii) Reflections. Achieving realistic rendering of reflective materials is a hard and long-standing problem in 3D scene reconstruction. GaussianShader [72] enhanced neural rendering for scenes with reflective surfaces by integrating a simplified shading function with 3D Gaussians. iii) Geometry. One limitation of 3D GS is the neglect of underlying scene geometry and structure, particularly in complex scenes and under varying view and lighting conditions. Scaffold-GS [73] introduced a sparse grid of anchor points to organize local 3D Gaussians, which are dynamically adjusted for attributes like opacity and color based on the viewer’s perspective and distance. This method enhanced scene representation through a hierarchical structure informed by scene geometry.
@@ 4.4 Improved Optimization Algorithms
Three main directions stand out for improving the optimization of 3D GS. i) Introducing Additional Regularization. 3D GS often faces challenges with over-reconstruction, where sparse, large 3D Gaussians cause blur and artifacts due to poor representation in high-variance regions. To address this, FreGS [89] introduced a progressive frequency regularization method, which refines Gaussian densification from a frequency perspective. In terms of the global geometry consistency, GeoGaussian [82] focused on preserving the geometry of non-textured regions like walls and furniture, which tend to degrade over time. ii) Improving Optimization Algorithm. In addressing the challenges of scale and structural inhomogeneity in 3D GS, which result in blurred boundaries and sparse point clouds in untextured areas, Feng et al. [90] introduced an advanced splitting algorithm. This algorithm decomposed an $N$-dimensional Gaussian into two, aiming to preserve essential mathematical characteristics (e.g., opacity, position, and covariance) and enhancing surface conformity and uniformity. iii) Relaxing Constraints in Optimization. Reliance on external tools/algorithms for can introduce errors and cap the system’s performance potential. For instance, structure-from-motion (SfM), commonly used in the initialization process, are error-prone and struggle with complex scenes. Yang et al. [86] proposed COLMAP-Free 3D GS, which introduces video stream continuity and an explicit point cloud representation so as to eliminate the need for SfM preprocessing.
@@
Here we list several interesting applications using 3D Gaussians with specially designed properties. i) Language Embedded Scene Representation. Due to the high computational and memory demands of current language-embedded scene representations, Shi et al. [91] proposed a quantization scheme that augments 3D Gaussian with streamlined language embeddings instead of the original high-dimensional embeddings. This method also mitigated semantic ambiguity and enhanced the precision of open-vocabulary querying by smoothing out semantic features across different views, guided by uncertainty values. ii) Scene Understanding and Editing. Feature 3DGS [94] integrated 3D GS with feature field distillation from 2D foundation models. By learning a lower-dimensional feature field and applying a lightweight convolutional decoder for upsampling, Feature 3DGS achieved faster training and rendering speeds while enabling high-quality feature field distillation, supporting applications like semantic segmentation and language-guided editing. iii) Spatiotemporal Modeling. To capture the complex spatial and temporal dynamics of 3D scenes, Yang et al. [97] conceptualized spacetime as a unified entity and approximates the spatiotemporal volume of dynamic scenes using a collection of 4D Gaussians. The proposed 4D Gaussian representation and corresponding rendering pipeline are capable of modeling arbitrary rotations in space and time and allow for end-to-end training.
@@
In addition to augmenting 3D Gaussian with additional properties, another promising avenue of adapting to downstream tasks is to introduce structured information (e.g., spatial MLPs and grids) tailored for specific applications.
Next we showcase various fascinating uses of 3D GS with specially devised structured information. i) Facial Expression Modeling. Considering the challenge of creating high-fidelity 3D head avatars under sparse view conditions, Gaussian Head Avatar [100] introduced controllable 3D Gaussians and an MLP-based deformation field. Concretely, it captured detailed facial expressions and dynamics by optimizing neutral 3D Gaussians alongside the deformation field, thus ensuring both detail fidelity and expression accuracy. ii) Spatiotemporal Modeling. Yang et al. [98] proposed to reconstruct dynamic scenes with deformable 3D Gaussians. The deformable 3D Gaussians are learned in a canonical space, coupled with a deformation field (i.e., a spatial MLP) that models the spatial-temporal dynamics. The proposed method also incorporated an annealing smoothing training mechanism to enhance temporal smoothness without additional computational costs. iii) Style Transfer. Saroha et al. [103] proposed GS in style, an advanced approach for real-time neural scene stylization. To maintain a cohesive stylized appearance across multiple views without compromising on rendering speed, they used pre-trained 3D Gaussians coupled with a multi-resolution hash grid and a small MLP to produce stylized views. Such a fast, view-consistent pipeline makes it particularly suited for AR/VR applications.
@@ 5 Application Areas and Tasks
The transformative potential of 3D GS extends far beyond its theoretical and computational advancements. This section delves into the diverse and groundbreaking application areas where 3D GS is making a significant impact, such as robotics (Sec. 5.1), dynamic scene reconstruction and representation (Sec. 5.2), AI-generated content (Sec. 5.3), autonomous driving (Sec. 5.4), medical systems (Sec. 5.5 and Sec. 5.6), and even other scientific disciplines. The applications of 3D GS demonstrate its versatility to revolutionize various fields. Here, we outline some of the most notable application domains, providing insights into how 3D GS is shaping new frontiers in each area.

### v3

@@ 3.2.1 Parameter Optimization
$\bullet$ Loss Function. Once the synthesis of the image is completed, the difference between the rendered image and ground truth can be measured. All the learnable parameters are optimized by stochastic gradient descent using the $\ell_{1}$ and D-SSIM loss functions:
@@
where $\lambda\in[0,1]$ is a weighting factor. The loss function of 3D GS is slightly different from that of NeRFs. NeRFs typically calculate the loss at the pixel level rather than the image level because of the costly ray-marching.
@@ 3.2.2 Density Control
$\bullet$ Initialization. 3D GS starts with the initial set of sparse points from SfM or random initialization. Note that a good initialization is essential to convergence and reconstruction quality [45]. Afterwards, point densification and pruning are adopted to control the density of 3D Gaussians.
$\bullet$ Point Densification. In the point densification phase, 3D GS adaptively increases the density of Gaussians to better capture the details of a scene. This process focuses on areas with missing geometric features or regions where Gaussians are too spread out. The densification procedure will be performed at regular intervals (i.e., after a certain number of training iterations), focusing on those Gaussians with large view-space positional gradients (i.e., above a specific threshold). It involves either cloning small Gaussians in under-reconstructed areas or splitting large Gaussians in over-reconstructed regions. For cloning, a copy of the Gaussian is created and moved towards the positional gradient. For splitting, a large Gaussian is replaced with two smaller ones, reducing their scale by a specific factor. This step seeks an optimal distribution and representation of Gaussians in 3D space, enhancing the overall quality of the reconstruction.
$\bullet$ Point Pruning. The point pruning stage involves the removal of superfluous or less impactful Gaussians, which can be viewed as a regularization process. It is executed by eliminating Gaussians that are virtually transparent (with $\alpha$ below a specified threshold) and those that are excessively large in either world-space or view-space. In addition, to prevent unjustified increases in Gaussian density near input cameras, the alpha value of the Gaussians is set close to zero after a certain number of iterations. This allows for a controlled increase in the density of necessary Gaussians while enabling the culling of redundant ones. The process not only helps in conserving computational resources but also ensures that the Gaussians in the model remain precise and effective for the representation of the scene.
@@
Though 3D GS has achieved impressive milestones, significant room for improvement remains, e.g., data and hardware requirement, rendering and optimization algorithm, and applications in downstream tasks. In the subsequent sections, we seek to elaborate on select extended versions of 3D GS. These are: i) Data-efficient 3D GS [46, 47, 48, 49, 50, 51, 52, 53, 54, 55] (Sec. 4.1), ii) Memory-efficient 3D GS [56, 57, 58, 59, 60, 61, 62, 63, 64] (Sec. 4.2), iii) Photorealistic 3D GS [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80] (Sec. 4.3), iv) Improved Optimization Algorithms [81, 22, 82, 77, 83, 84, 85, 86] (Sec. 4.4), v) 3D Gaussian with More Properties [87, 88, 89, 90, 91, 92, 93] (Sec. 4.5), and vi) 3D GS with Structured Information [94, 95, 96] (Sec. 4.6).
@@ 4.1 Data-efficient 3D GS
A notable issue of 3D GS is the emergence of artifacts in areas with insufficient observational data. This challenge is a prevalent limitation in radiance field rendering, where sparse data often leads to inaccuracies in reconstruction. From a practical perspective, reconstructing scenes from limited viewpoints is of significant interest, particularly for the potential to enhance functionality with minimal input.
Two primary strategies are used for data-efficient 3D GS. i) Regularization based methods introduce additional constraints such as depth information to enhance the detail and global consistency [47, 50, 52]. For example, DNGaussian [50] introduced a depth-regularized approach to address the challenge of geometry degradation in sparse input views through hard and soft depth regularization and global-local depth normalization. It rectified the geometry using depth information extracted from pre-trained monocular depth estimators and optimized the positioning of Gaussian primitives without changing their shape. Unfortunately, when dealing with a limited number of views, or even just one, the efficacy of regularization techniques tends to diminish, which leads to ii) generalizability based methods that focus in particular on learning priors [48, 49, 54]. One typical implementation is to generate 3D Gaussians that can be used for rendering directly without optimization, by using deep neural networks. This paradigm typically requires multiple views for training but can reconstruct the 3D scene with only one input image. For instance, PixelSplat [48] proposed to sample Gaussians from dense probability distributions. It incorporated a multi-view epipolar transformer and a reparameterization trick to avoid local minima and maintain gradient flow. Splatter Image [49] applied GS in a monocular setting through a learning-based approach, utilizing a 2D image-to-image network that maps an input image to a 3D Gaussian per pixel. Note that this paradigm is mainly focused on the reconstruction of objects, and its generalizability leaves a lot of room for improvement.
@@ 4.2 Memory-efficient 3D GS
While 3D GS demonstrates remarkable capabilities, its scalability poses significant challenges, particularly when juxtaposed with NeRF-based methods. The latter benefits from the simplicity of storing merely the parameters of a learned MLP. This scalability issue becomes increasingly acute in the context of large-scale scene management, where the computational and memory demands escalate substantially. Consequently, there is an urgent need to optimize memory usage in both model training and storage.
There are two main directions to reduce memory usage. i) The first involves reducing the number of 3D Gaussians [63, 58, 62], i.e., pruning insignificant 3D Gaussians. For example, Papantonakis et al. [63] proposed a resolution-aware pruning method to prune Gaussians, reducing the Gaussian count by half. Lee et al. [58] introduced a new volume-based masking strategy that efficiently reduces the quantity of Gaussians without compromising performance. ii) The second category focuses on compressing the memory usage of 3D Gaussian properties [61, 58, 62]. For instance, Niedermayr et al. [61] compressed color and Gaussian parameters into compact codebooks, using sensitivity measures for effective quantization and fine-tuning. HAC [62] predicted the probability of each quantized attribute using Gaussian distributions and then devise an adaptive quantization module. Although current methods achieve compression ratios of several to dozens of times for storage (after training), there remains considerable potential for reducing memory usage during the training phase.
@@ 4.3 Photorealistic 3D GS
The current rendering pipeline of 3D GS (Sec. 3.1) is straightforward and involves several drawbacks. For instance, the simple visibility algorithm may lead to a drastic switch in the depth/blending order of Gaussians [10]. The realness of rendered images, including aspects such as aliasing, reflections, and artifacts, can be further optimized.
Here we list several key points for enhancing realness. i) Varying Resolutions [67, 78]. Due to the discrete sampling paradigm (viewing each pixel as a single point instead of an area), 3D GS is susceptible to aliasing when dealing with varying resolutions, which leads to blurring or jagged edges. Yan et al. [67] argued that this is mainly because traditional rendering methods cannot effectively manage the disparity between the sampling frequency of pixels and the high-frequency details of the scene, leading to visual artifacts and performance issues. Hence, they introduced multi-scale 3D GS, where the scene is represented using Gaussians of varying sizes. Analytic-Splatting [78] adopted an analytic approximation of the Gaussian integral within the pixel area, leveraging a conditioned logistic function for the cumulative distribution function to better capture the intensity response of pixels. ii) Reflections [68, 97, 98]. Achieving realistic rendering of reflective materials is a hard and long-standing problem in 3D scene reconstruction. GaussianShader [68] enhanced neural rendering for scenes with reflective surfaces by integrating a simplified shading function with 3D Gaussians. iii) Geometry. One limitation of 3D GS is the neglect of underlying scene geometry and structure, particularly in complex scenes and under varying view and lighting conditions. This sparks research on geometry-aware reconstruction [99, 100, 45, 22, 101, 102]. For example, GeoGaussian [77] focused on preserving the geometry of non-textured regions like walls and furniture, which tend to degrade over time.
@@ 4.4 Improved Optimization Algorithms
Three main directions stand out for improving the optimization of 3D GS. i) Introducing Additional Regularization [84, 22]. 3D GS often faces challenges with over-reconstruction, where sparse, large 3D Gaussians cause blur and artifacts due to poor representation in high-variance regions. To address this, FreGS [84] introduced a progressive frequency regularization method, which refines Gaussian densification from a frequency perspective. Another notable branch is geometry-aware reconstruction, as introduced in Sec. 4.3. This line of work focuses in particular on how to preserve scene’s structure. For instance, Scaffold-GS [22] introduced a sparse grid of anchor points to organize local 3D Gaussians, which are dynamically adjusted for attributes like opacity and color based on the viewer’s perspective and distance. ii) Improving Optimization Procedure [45, 77]. In addressing the challenges of dense initialization for texture-less surfaces, especially in large-scale scenes, GaussianPro [45] devised an advanced Gaussian densification (§3.2.2) strategy using the priors of reconstructed geometries and patch matching techniques. iii) Relaxing Constraints in Optimization [81, 82]. Reliance on external tools/algorithms for can introduce errors and cap the system’s performance potential. For instance, SfM, commonly used in the initialization process, is error-prone and struggle with complex scenes. Yang et al. [81] proposed COLMAP-Free 3D GS, which introduces video stream continuity and an explicit point cloud representation so as to eliminate the need for SfM preprocessing. Though impressive, existing methods primarily concentrate on optimizing Gaussians to accurately reconstruct scenes from scratch, neglecting a challenging yet promising paradigm which reconstructs scenes in a few-shot manner through established “meta representations”. See “learning physical priors from large-scale data” in Sec. 7 for further insights.
@@
Here we list several interesting applications using 3D Gaussians with specially designed properties. i) Language Embedded Scene Representation [87, 88, 89]. Due to the high computational and memory demands of current language-embedded scene representations, Shi et al. [87] proposed a quantization scheme that augments 3D Gaussian with streamlined language embeddings instead of the original high-dimensional embeddings. This method also mitigated semantic ambiguity and enhanced the precision of open-vocabulary querying by smoothing out semantic features across different views, guided by uncertainty values. ii) Scene Understanding and Editing [90, 91, 92]. Feature 3DGS [90] integrated 3D GS with feature field distillation from 2D foundation models. By learning a lower-dimensional feature field and applying a lightweight convolutional decoder for upsampling, Feature 3DGS achieved faster training and rendering speeds while enabling high-quality feature field distillation, supporting applications like semantic segmentation and language-guided editing. iii) Spatiotemporal Modeling [93, 103]. To capture the complex spatial and temporal dynamics of 3D scenes, Yang et al. [93] conceptualized spacetime as a unified entity and approximates the spatiotemporal volume of dynamic scenes using a collection of 4D Gaussians. The proposed 4D Gaussian representation and corresponding rendering pipeline are capable of modeling arbitrary rotations in space and time and allow for end-to-end training.
[104], [105], [106], [107], [108], [109], [110],[111], [112], [113], [114], [115], [116], [117][118], [94], [95], [93], [119], [120],[121], [122], [103], [123], [124], [125][126], [127], [128], [129], [130], [131],[132], [133], [134], [135], [90], [136][137], [138], [139][140], [141], [142][143], [116], [144][45], [145], [146], [147], [148], [149], [150], [151]
Figure 6: Typical applications that have been benefited from GS (Sec. 5). Some images are borrowed from [125, 128, 138, 142, 151] and redrawn.
@@
Rather than augmenting 3D Gaussian with additional properties, another promising avenue of adapting to downstream tasks is to introduce structured information (e.g., spatial MLPs and grids) tailored for specific applications.
Next we showcase various fascinating uses of 3D GS with specially devised structured information. i) Facial Expression Modeling. Considering the challenge of creating high-fidelity 3D head avatars under sparse view conditions, Gaussian Head Avatar [96] introduced controllable 3D Gaussians and an MLP-based deformation field. Concretely, it captured detailed facial expressions and dynamics by optimizing neutral 3D Gaussians alongside the deformation field, thus ensuring both detail fidelity and expression accuracy. ii) Spatiotemporal Modeling. Yang et al. [94] proposed to reconstruct dynamic scenes with deformable 3D Gaussians. The deformable 3D Gaussians are learned in a canonical space, coupled with a deformation field (i.e., a spatial MLP) that models the spatial-temporal dynamics. The proposed method also incorporated an annealing smoothing training mechanism to enhance temporal smoothness without additional computational costs. iii) Style Transfer. Saroha et al. [152] proposed GS in style, an advanced approach for real-time neural scene stylization. To maintain a cohesive stylized appearance across multiple views without compromising on rendering speed, they used pre-trained 3D Gaussians coupled with a multi-resolution hash grid and a small MLP to produce stylized views. In a nutshell, incorporating structured information can serve as a complementary part for adapting to tasks that are incompatible with the sparsity and disorder of 3D Gaussians.
@@ 5 Application Areas and Tasks
Building on the rapid advancements in 3D GS, a wide range of innovative applications has emerged across multiple domains (Fig. 6) such as robotics (Sec. 5.1), dynamic scene reconstruction and representation (Sec. 5.2), AI-generated content (Sec. 5.3), autonomous driving (Sec. 5.4), medical systems (Sec. 5.5), large-scale scene reconstruction (Sec. 5.6), and even other scientific disciplines [153, 154, 24, 155]. Here, we highlight key examples that underscore the transformative impact and potential of 3D GS.